# Agnostic Process Tomography

### v1

@@ 5.3 Pauli strings and Pauli channels
$\hat{\Phi}(x,x)=p_{x},\quad\forall x\in\mathcal{P}_{n}$ (5.64)
@@
$|\hat{p}(x)-\hat{\Phi}(x,x)|\leq\epsilon_{1},\quad\forall x\in\mathcal{P}_{n},$ (5.77)
@@
$\hat{\Phi}(x^{*},x^{*})+\epsilon_{1}\geq\hat{p}(x^{*})\quad\text{ and }\quad\hat{p}(z^{\star})\geq\hat{\Phi}(z^{*},z^{*})-\epsilon_{1}.$ (5.79)
@@
1:Prepare $N=\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ copies of $v(\Phi)$ and measure them in the {|$\{\ket{v(\sigma_{x})}\}_{x}$ basis to construct estimates $\hat{p}(x)$ of $\hat{\Phi}(x,x)$.
@@ 5.4 Junta channels
However, in the case of a quantum channel $\Phi$, a similar method allows us to sample from the distribution $\{\hat{\Phi}(x,x)\}_{x\in\mathcal{P}_{n}}$. This does not allow us to estimate the weight of a subset $S$, which is given by $w_{S}(\Phi)=\sum_{x,y\in\mathcal{P}_{S}}|\hat{\Phi}(x,y)|^{2}$. To overcome this issue, we learn all Fourier coefficients of degree at most $k$, and use them to estimate the weight. Unsurprisingly, this results in a worse sample complexity than the unitary junta case (Theorem 2) Once we have estimated the weights, we can construct an improper hypothesis by restricting to the Fourier coefficients corresponding to the heaviest estimated subset. The algorithm is sketched in Algorithm 6.
@@
1:Use $N$ queries to $\Phi$ to construct estimates $\alpha_{x,y}$ of the Fourier coefficients for all $x,y\in\mathcal{P}_{n}$ such that $|x|,|y|\leq k$.
@@
Now that we have lower bounded opt, we analyze our algorithm (Algorithm 6), which starts by estimating all Fourier coefficients of degree at most $k$, Note that there are at most $n^{k}16^{k}$ Fourier coefficients of degree at most $k$. Thus, we use Lemma 12 $\mathcal{O}(n^{k}16^{k})$ times, to obtain estimates $\alpha_{x,y}$, such that with probability at least $1-\delta$,
$|\alpha_{x,y}-\hat{\Phi}(x,y)|\leq\epsilon_{1},\quad\forall x,y\in\mathcal{P}_{n}\text{ s.t. }|x|,|y|\leq k.$ (5.109)
@@
$\hat{w}_{S}\triangleq\sum_{x,y\in\mathcal{P}_{S}}|\alpha_{x,y}|^{2},\quad\forall S\subseteq[n],|S|=k.$ (5.111)
Note that all Fourier coefficients in these summations have a degree of at most $k$. Thus, by the previous step of estimating all Fourier coefficients of degree at most $k$ by $\alpha_{x,y}$, we can compute $\hat{w}_{S}$. As there are $\mathcal{O}(n^{k})$ subsets of cardinality $k$, and computing each weight requires adding weights of $16^{k}$ coefficients, finding all the weights takes time $\mathcal{O}(n^{k}\cdot 16^{k})$. We now bound the error of these estimates.
@@
$w_{S^{\prime}}(\Phi)+2\cdot 16^{k}\epsilon_{1}\geq\hat{w}_{S^{\prime}}\quad\text{ and }\quad\hat{w}_{S^{\star}}\geq w_{S^{\star}}(\Phi)-2\cdot 16^{k}\epsilon_{1}.$ (5.127)
@@ 5.5 Low-degree channels
We note a difference between our definition of degree-$d$ channels and that of [30]. Our definition requires that for all $\hat{\Phi}(x,y)\neq 0$, both $|x|$ and $|y|$ are at most $d$. On the other hand, the definition of [30] requires that for all non-zero $\hat{\Phi}(x,y)$, $|x|+|y|\leq d$.
@@
1:Use $N$ queries to $\Phi$ to construct estimates $\alpha_{x,y}$ of the Fourier coefficients for all $x,y\in\mathcal{P}_{n}$ such that $|x|,|y|\leq d$.
@@
$N=\mathcal{O}\left(\frac{d\cdot 256^{d}n^{2d}\log(n/\delta)}{\epsilon^{2}}\right)$ (5.133)
@@
Our algorithm (Algorithm 7) estimates all Fourier coefficients of degree at most $d$. Note that there are at most $n^{d}16^{d}$ of them. We use Lemma 12 $\mathcal{O}(n^{d}16^{d})$ times, to obtain estimates $\alpha_{x,y}$ such that
$|\alpha_{x,y}-\hat{\Phi}(x,y)|\leq\epsilon_{1},\quad\forall x,y\in\mathcal{P}_{n}\text{ s.t. }|x|,|y|\leq d,$ (5.140)
@@
$N=\mathcal{O}\left(\frac{d\cdot 16^{d}n^{d}\log(n/\delta)}{\epsilon_{1}^{2}}\right).$ (5.141)
@@
$\leq\frac{n^{d}16^{d}\epsilon_{1}^{2}}{2}.$ (5.151)
In the second line, we use the definition of $\Phi_{d}$. In the third line, we use Equation 5.140 and that there are at most $n^{d}16^{d}$ Fourier coefficients of degree at most $d$. Thus, with probability at least $1-\delta$, we have
$d_{F}(\Phi,\Psi)\leq\mathsf{opt}(\Phi,\mathcal{LC}_{n,d})+\frac{\sqrt{n^{d}16^{d}}\epsilon_{1}}{\sqrt{2}}.$ (5.152)
To obtain the 1-agnostic learning guarantee, we choose $\epsilon_{1}=\epsilon\sqrt{2}/\sqrt{n^{d}16^{d}}$. This gives us the stated query and time complexities. ∎
@@ 5.6 $n$-to-1 QAC0 channels
In this section, we consider the class of channels $\mathsf{QAC}^{0}$ formed by constant-depth circuits with arbitrary single-qubit gates and arbitrary-width phase-flip gates [86]. Following standard terminology, we refer to the size of a quantum circuit as the number of multi-qubit gates, and the depth as the number of layers of multi-qubit gates in the circuit. We consider the algorithm of [22] and show that it is an improper 1-agnostic learning algorithm. We follow the definition of [22] for the circuit class $\mathsf{QAC}^{0}$.
Definition 20 ($\mathsf{QAC}^{0}$ circuits; Definition 16 in [22]).
The class $\mathsf{QAC}^{0}$ contains quantum circuit families consisting of $\mathsf{CZ}$ gates and arbitrary single-qubit gates, with depth bounded below a constant. That is, the circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ is in $\mathsf{QAC}^{0}$ if there exists a $d\in\mathbb{N}$ such that for each $n\in\mathbb{N}$, $C_{n}$ is a quantum circuit on $n$ qubits consisting only of $\mathsf{CZ}$ gates and arbitrary single-qubit gates, with depth at most $d$.
The algorithm of [22] focuses on learning a specific class of channels formed by $\mathsf{QAC}^{0}$ circuits. In particular, they consider the class of channels formed by tracing out all except a single qubit.
Definition 21 ($n$-to-1 $\mathsf{QAC}^{0}$ channels; Definition 12 in [22]).
Let $n,a\in\mathbb{N}$. Let $U$ be an $(n+a)$-qubit unitary implemented by a $\mathsf{QAC}^{0}$ circuit, and let $\psi$ be an $a$-qubit quantum state. We define the $n$-to-$1$ $\mathsf{QAC}^{0}$ channel $\mathcal{E}_{U,\psi}$ to be the following channel mapping $n$ qubits to $1$ qubit
@@
[…]
We consider the Pauli spectrum expansion (see Definition 4) of the Choi representation of such channels. Namely, for an $n$-to-$1$ $\mathsf{QAC}^{0}$ channel $\mathcal{E}_{U,\psi}$, the Choi representation $J(\mathcal{E}_{U,\psi})\in\mathcal{B}((\mathbb{C}^{2})^{\otimes n+1})$ can be expressed as
@@
One of the key technical results of [22] is that when expressing the Choi representation of $n$-to-1 $\mathsf{QAC}^{0}$ channels in the Pauli spectrum, the weight is concentrated on the low-degree terms. Formally, we have the following lemma.
Lemma 18 (Low-degree concentration of $\mathsf{QAC}^{0}$ channels; Theorem 18 in [22]).
@@
Algorithm 8 1-Agnostic Improper Learning $n$-to-$1$ $\mathsf{QAC}^{0}$ Channels
1:Prepare $N$ copies of $v(\Phi)$.
@@
$\mathcal{O}_{k}\triangleq\{A\in\mathcal{B}((\mathbb{C}^{2})^{\otimes n+1}):w^{>k}(A)=0\}.$ (5.158)
@@
$|\hat{\beta}_{x}-2\alpha_{x}|\leq\epsilon_{1},\quad\forall x\in\mathcal{P}_{n+1}\text{ s.t. }|x|\leq k$ (5.176)
@@
Unlike the case for APT of junta channels and low-degree channels, the method discussed in Section 5.7 does not help us obtain a proper hypothesis for $\mathsf{QAC}^{0}$ channels. As discussed in [22], one can map the problem of obtaining the nearest quantum channel from the improper hypothesis to a convex optimization problem. However, it is not clear how to obtain the nearest quantum channel that can be implemented using a $\mathsf{QAC}^{0}$ circuit. As a result, we only achieve improper agnostic process tomography of $\mathsf{QAC}^{0}$ channels.
@@ 5.7 Proper APT
In this section, we discuss an approach towards proper learning for APT based on convex optimization. In particular, we aim to devise a projection algorithm which takes as input an improper hypothesis output by one of our algorithms (e.g., Sections 5.4 and 5.5) and returns a valid quantum channel. We are unable to fully solve this problem due to a difficult-to-verify condition from [102], which we rephrase in our setting in as Conjecture 1. Nonetheless, we consider this as a significant first step towards resolving this open question, which is also recurrent in much of the realizable process tomography literature that uses Pauli spectrum analysis [22, 24, 25]. Note that ideas from the quasiprobability decomposition and probabilistic error cancellation which approximate linear superoperators by linear combinations of quantum channels [103, 104, 105, 81, 106, 107, 108, 109, 110] do not help because we need to find exactly a quantum channel.
Suppose that we are given an improper hypothesis $\Psi_{1}$, from an agnostic process tomography algorithm of a concept class $\mathcal{C}$ and an unknown channel $\Phi$, with respect to the normalized Frobenius distance $d_{F}$. Denote the Fourier coefficient matrix of $\Phi$ by $\hat{\Phi}$. Moreover, suppose that our hypothesis is given to us as a Hermitian matrix of Fourier coefficients, $\hat{\Psi}_{1}$. This aligns with our learning algorithms from Sections 5.4 and 5.5, where we learn the Fourier coefficients corresponding to a principle submatrix of $\hat{\Phi}$ and simply set the rest to 00. To ensure that $\hat{\Psi}_{1}$ is Hermitian, we can instead only learn the upper triangular coefficients of this principle submatrix, and set the corresponding coefficients in the lower triangle as their complex conjugates. From the improper APT guarantee, we have
@@
[…]
@@
This transformation is well-known, first noted in [146], and can be used to transform any complex semidefinite programming problem into its real analogue. It can be shown that this transformation preserves the properties of the original operator. For example, a Hermitian operator is mapped to a symmetric one and a Hermitian positive semi-definite operator is mapped to a symmetric positive semi-definite one. Using this transformation, we see that it suffices to solve the following convex optimization problem over reals:
@@
$C_{1}=\{x\in\mathbb{R}^{4D^{2}}:\mathsf{vec}^{-1}(x)\in S_{2D}^{+}\},\quad C_{2}=\{x\in\mathbb{R}^{4D^{2}}:\tilde{A}x=\tilde{b}\},$ (5.211)
@@
Note that projecting onto either $C_{1}$ or $C_{2}$ individually is easy (running in time $\mathrm{poly}(D)$) since they are simply the PSD cone and a polyhedral set (see, e.g., [147]). Note that the runtime being in terms of $D$ is not a problem for our applications, as we discuss at the end of this section. Thus, the problem is as follows: how can we use projections onto $C_{1}$ and $C_{2}$ individually to project onto $C_{1}\cap C_{2}$? This problem is well-studied in convex optimization, where researchers have developed methods such as alternating projections [148, 149] and Dykstra’s algorithm [150]. As the name suggests, at a high level, these algorithms work by projecting the input onto $C_{1}$ and $C_{2}$ in turn. Classical results in optimization theory show that Dykstra’s algorithm converges to the desired projection [150, 151, 152, 153, 154], and its variants are also widely studied [155, 156, 157, 158, 159]; however, researchers have only recently rigorously analyzed its convergence rate for non-polyhedral sets [102]. Convergence rates have previously been analyzed for the basic alternating projections algorithm [160, 161, 162] and Dykstra’s algorithm for projection onto the intersection of polyhedral sets [156, 163, 164]. However, the former only guarantees convergence to a point in $C_{1}\cap C_{2}$, rather than the closest one, which would not help for our purposes since we require an error bound. Also, the latter does not apply to our setting, as $C_{1}=S_{2D}^{+}$ is clearly not polyhedral. In our setting, using Dykstra’s algorithm, one would hope for linear convergence, as this leads to a computationally efficient (in the size of the input matrix) projection algorithm. In fact, [102] shows that Dykstra’s algorithm achieves a linear convergence rate under certain conditions. Namely, they prove the following theorem.
Theorem 10 (Theorem 5.3(i) in [102]; Simplified).
@@
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\norm{x-x_{0}}_{2}^{2}\quad\text{s.t.}\quad x\in\bigcap_{i=1}^{\ell}C_{i},$ (5.212)
@@
    $0\in x^{*}-x_{0}+\mathrm{ri}\left(\partial\left(\sum_{i=1}^{\ell}\delta_{C_{i}}\right)(x^{*})\right)$, where $x^{*}$ is the optimal projected point, $\partial$ denotes the subdifferential, and $\delta_{C_{i}}$ is the indicator function of the set $C_{i}$ (00 for inputs inside $C_{i}$ and +$\infty$+ ∞ otherwise).
@@
Note that [102] considers a more general problem, in which they wish to project onto a point such that a linear transformation of this point lies in the intersection of the convex sets. For our setting, we can take this transformation to be the identity, in which case the algorithm from [102] reduces to Dykstra’s algorithm. Thus, this theorem would give the convergence result we need if its conditions hold.
Now, let us discuss each of the conditions. First, for Condition 1, the exact definition of $C^{1,\alpha}$-cone reducibility, for $\alpha\in(0,1]$, is outside of the scope of this work, and we refer to Definition 4.1 in [102] for the definition. Importantly, this is an extension of the more common notion of $C^{2}$-cone reducibility [165, 166]. Specifically,  [102] states that any $C^{2}$-cone reducible set is also $C^{1,\alpha}$-cone reducible for all $\alpha\in(0,1]$. Moreover, [165] shows that all polyhedrons and the cone of positive semidefinite matrices are both $C^{2}$-cone reducible. Hence, our convex sets $C_{1}$ and $C_{2}$ are both $C^{2}$-cone reducible and hence $C^{1,\alpha}$-cone reducible for all $\alpha\in(0,1]$. In particular, $C_{1}$ and $C_{2}$ are both $C^{1,1}$-cone reducible, satisfying Condition 1.
@@
Also, $\partial$ denotes the subdifferential. For a function $f:\mathbb{R}^{4D^{2}}\to\mathbb{R}$, the subdifferential is defined as
$\partial f(x)\triangleq\{z\in\mathbb{R}^{4D^{2}}:f(y)\geq f(x)+(y-x)^{\intercal}z\quad\forall y\in\mathbb{R}^{4D^{2}}\}.$ (5.214)
@@
$\mathrm{ri}(S)\triangleq\{x\in S:\forall y\in S,\;\exists\lambda>1\text{ s.t. }\lambda x+(1-\lambda)y\in S\}.$ (5.215)
Because of Condition 2, standard results in subdifferential calculus (see, e.g., [167]) state that
@@
$N_{S}(y)\triangleq\{a\in\mathbb{R}^{4D^{2}}:a^{\intercal}y\geq a^{\intercal}x,\quad\forall x\in S\}.$ (5.217)

### v2

@@ 5.3 Pauli strings and Pauli channels
$\hat{\Phi}(x,x)=p_{x},\qquad\forall x\in\mathcal{P}_{n}$ (5.64)
@@
$|\hat{p}(x)-\hat{\Phi}(x,x)|\leq\epsilon_{1},\qquad\forall x\in\mathcal{P}_{n},$ (5.77)
@@
$\hat{\Phi}(x^{*},x^{*})+\epsilon_{1}\geq\hat{p}(x^{*})\qquad\text{ and }\qquad\hat{p}(z^{\star})\geq\hat{\Phi}(z^{*},z^{*})-\epsilon_{1}.$ (5.79)
@@
[…]
@@ 5.4 Junta channels
However, in the case of a quantum channel $\Phi$, a similar method allows us to sample from the distribution $\{\hat{\Phi}(x,x)\}_{x\in\mathcal{P}_{n}}$. This does not allow us to estimate the weight of a subset $S$, which is given by $w_{S}(\Phi)=\sum_{x,y\in\mathcal{P}_{S}}|\hat{\Phi}(x,y)|^{2}$. To overcome this issue, we learn the Fourier coefficients associated with each subset and use them to estimate the weight. Unsurprisingly, this results in a worse sample complexity than the unitary junta case (Theorem 2). Once we have estimated the weights, we can construct an improper hypothesis by restricting to the Fourier coefficients corresponding to the heaviest estimated subset. The algorithm is sketched in Algorithm 6.
@@
1:Use NN queries to $\Phi$ to construct estimates $\alpha_{x,y}$ of the Fourier coefficients with $x,y\in\mathcal{P}_{S}$, for all subsets $S\subseteq[n],|S|=k$.
@@
Now that we have lower bounded opt, we analyze our algorithm (Algorithm 6), which starts by estimating all Fourier coefficients associated with subsets of size $k$. Note that there are at most $n^{k}16^{k}$ such Fourier coefficients. Thus, we use Lemma 12 $\mathcal{O}(n^{k}16^{k})$ times, to obtain estimates $\alpha_{x,y}$, such that with probability at least $1-\delta$,
$|\alpha_{x,y}-\hat{\Phi}(x,y)|\leq\epsilon_{1},\qquad\forall x,y\in\mathcal{P}_{S},\;S\subseteq[n],\;|S|=k.$ (5.109)
@@
$\hat{w}_{S}\triangleq\sum_{x,y\in\mathcal{P}_{S}}|\alpha_{x,y}|^{2},\qquad\forall S\subseteq[n],|S|=k.$ (5.111)
As we have already estimated all Fourier coefficients associated with subsets of cardinality $k$ by $\alpha_{x,y}$, we can now compute $\hat{w}_{S}$. As there are $\mathcal{O}(n^{k})$ subsets of cardinality $k$, and computing each weight requires adding weights of $16^{k}$ coefficients, finding all the weights takes time $\mathcal{O}(n^{k}\cdot 16^{k})$. We now bound the error of these estimates.
@@
[…]
@@ 5.5 Low-degree channels
We note a difference between our definition of degree-$d$ channels and that of [arunachalam2024learning]. Our definition requires that for all $\hat{\Phi}(x,y)\neq 0$, both $|x|$ and $|y|$ are at most $d$. On the other hand, the definition of [arunachalam2024learning] requires that for all non-zero $\hat{\Phi}(x,y)$, $|x|+|y|\leq d$. With our chosen definition, the channels corresponding to degree-$d$ unitaries (Definition 16) are also channels of degree-$d$. Thus, we find this to be a more natural definition, and we also use this fact in Remark 3. However, the definitions do not differ significantly, and one can show that an approach similar to Algorithm 7 allows for agnostic tomography of low-degree channels adhering to the definition in [arunachalam2024learning].
@@
1:Use NN queries to $\Phi$ to construct estimates $\alpha_{x,y}$ of the Fourier coefficients for all $x,y\in\mathcal{P}_{n}$ such that $|x|,|y|\leq d$.
@@
$N=\mathcal{O}\left(\frac{d\cdot 81^{d}n^{4d}\log(n/\delta)}{\epsilon^{2}}\right)$ (5.133)
@@
Our algorithm (Algorithm 7) estimates all Fourier coefficients of degree at most $d$. Note that there are at most $\mathcal{O}(n^{2d}9^{d})$ of them. We use Lemma 12 $\mathcal{O}(n^{2d}9^{d})$ times, to obtain estimates $\alpha_{x,y}$ such that
$|\alpha_{x,y}-\hat{\Phi}(x,y)|\leq\epsilon_{1},\qquad\forall x,y\in\mathcal{P}_{n}\text{ s.t. }|x|,|y|\leq d,$ (5.140)
@@
$N=\mathcal{O}\left(\frac{d\cdot 9^{d}n^{2d}\log(n/\delta)}{\epsilon_{1}^{2}}\right).$ (5.141)
@@
$\leq\mathcal{O}\left(n^{2d}9^{d}\epsilon_{1}^{2}\right).$ (5.151)
In the second line, we use the definition of $\Phi_{d}$. In the third line, we use Equation 5.140 and that there are at most $\mathcal{O}(n^{2d}9^{d})$ Fourier coefficients of degree at most $d$. Thus, with probability at least $1-\delta$, we have
$d_{F}(\Phi,\Psi)\leq\mathsf{opt}(\Phi,\mathcal{LC}_{n,d})+\mathcal{O}\left(n^{d}3^{d}\epsilon_{1}\right).$ (5.152)
To obtain the 1-agnostic learning guarantee, we choose $\epsilon_{1}=\mathcal{O}(\epsilon/n^{d}3^{d})$. This gives us the stated query and time complexities. ∎
@@ 5.6 $n$-to-1 QAC0 channels
In this section, we consider the class of channels $\mathsf{QAC}^{0}$ formed by constant-depth circuits with arbitrary single-qubit gates and controlled phase-flip gates of arbitrary width [moore1999quantum]. Following standard terminology, we refer to the size of a quantum circuit as the number of multi-qubit gates, and the depth as the number of layers of multi-qubit gates in the circuit. We consider the algorithm of [nadimpalli2024pauli] and show that it is an improper 1-agnostic learning algorithm. We follow the definition of [nadimpalli2024pauli] for the circuit class $\mathsf{QAC}^{0}$.
Definition 20 ($\mathsf{QAC}^{0}$ circuits; Definition 16 in [nadimpalli2024pauli]).
The class $\mathsf{QAC}^{0}$ contains quantum circuit families consisting of multiple-qubit $\mathsf{CZ}$ gates and arbitrary single-qubit gates, with depth bounded below a constant. That is, the circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ is in $\mathsf{QAC}^{0}$ if there exists a $d\in\mathbb{N}$ such that for each $n\in\mathbb{N}$, $C_{n}$ is a quantum circuit on $n$ qubits consisting only of multiple-qubit $\mathsf{CZ}$ gates and arbitrary single-qubit gates, with depth at most $d$.
The algorithm of [nadimpalli2024pauli] focuses on learning a specific class of channels formed by $\mathsf{QAC}^{0}$ circuits. In particular, they consider the class of channels formed by tracing out all except a single qubit.
Definition 21 ($n$-to-1 $\mathsf{QAC}^{0}$ channels; Definition 12 in [nadimpalli2024pauli]).
Let $n,a\in\mathbb{N}$. Let $U$ be an $(n+a)$-qubit unitary implemented by a $\mathsf{QAC}^{0}$ circuit, and let $\psi$ be an $a$-qubit quantum state. We define the $n$-to-11 $\mathsf{QAC}^{0}$ channel $\mathcal{E}_{U,\psi}$ to be the following channel mapping $n$ qubits to $1$ qubit
@@
where tr[$\tr_{[n+a-1]}$ denotes the partial trace over all except the last qubit.
We consider the Pauli spectrum expansion (see Definition 4) of the Choi representation of such channels. Namely, for an $n$-to-11 $\mathsf{QAC}^{0}$ channel $\mathcal{E}_{U,\psi}$, the Choi representation $J(\mathcal{E}_{U,\psi})\in\mathcal{B}((\mathbb{C}^{2})^{\otimes n+1})$ can be expressed as
@@
One of the key technical results of [nadimpalli2024pauli] is that when expressing the Choi representation of $n$-to-1 $\mathsf{QAC}^{0}$ channels in the Pauli spectrum, the weight is concentrated on the low-degree terms. Formally, we have the following lemma.
Lemma 18 (Low-degree concentration of $\mathsf{QAC}^{0}$ channels; Theorem 18 in [nadimpalli2024pauli]).
@@
Algorithm 8 1-Agnostic Improper Learning $n$-to-11 $\mathsf{QAC}^{0}$ Channels
1:Prepare NN copies of $v(\Phi)$.
@@
$\mathcal{O}_{k}\triangleq\{A\in\mathcal{B}((\mathbb{C}^{2})^{\otimes n+1})\mathrel{\mathop{\ordinarycolon}}w^{>k}(A)=0\}.$ (5.158)
@@
$|\hat{\beta}_{x}-2\alpha_{x}|\leq\epsilon_{1},\qquad\forall x\in\mathcal{P}_{n+1}\text{ s.t. }|x|\leq k$ (5.176)
@@
Unlike the case for APT of junta channels and low-degree channels, the method discussed in Section 5.7 does not help us obtain a proper hypothesis for $\mathsf{QAC}^{0}$ channels. As discussed in [nadimpalli2024pauli], one can map the problem of obtaining the nearest quantum channel from the improper hypothesis to a convex optimization problem. However, it is not clear how to obtain the nearest quantum channel that can be implemented using a $\mathsf{QAC}^{0}$ circuit. As a result, we only achieve improper agnostic process tomography of $\mathsf{QAC}^{0}$ channels.
@@ 5.7 Proper APT
In this section, we discuss an approach towards proper learning for APT based on convex optimization. In particular, we aim to devise a projection algorithm which takes as input an improper hypothesis output by one of our algorithms (e.g., Sections 5.4 and 5.5) and returns a valid quantum channel. We are unable to fully solve this problem due to a difficult-to-verify condition from [wang2024convergence], which we rephrase in our setting in as Conjecture 1. Nonetheless, we consider this as a significant first step towards resolving this open question, which is also recurrent in much of the realizable process tomography literature that uses Pauli spectrum analysis [nadimpalli2024pauli, chen2023testing, bao2023testing]. Note that ideas from the quasiprobability decomposition and probabilistic error cancellation which approximate linear superoperators by linear combinations of quantum channels [jiang2021physical, regula2021operational, temme2017error, van2023probabilistic, zhao2023power, horodecki2002method, horodecki2003limits, fiuravsek2002structural, korbicz2008structural] do not help because we need to find exactly a quantum channel.
Suppose that we are given an improper hypothesis $\Psi_{1}$, from an agnostic process tomography algorithm of a concept class $\mathcal{C}$ and an unknown channel $\Phi$, with respect to the normalized Frobenius distance $d_{F}$. Denote the Fourier coefficient matrix of $\Phi$ by $\hat{\Phi}$. Moreover, suppose that our hypothesis is given to us as a Hermitian matrix of Fourier coefficients, $\hat{\Psi}_{1}$. This aligns with our learning algorithms from Sections 5.4 and 5.5, where we learn the Fourier coefficients corresponding to a principle submatrix of $\hat{\Phi}$ and simply set the rest to 0. To ensure that $\hat{\Psi}_{1}$ is Hermitian, we can instead only learn the upper triangular coefficients of this principle submatrix, and set the corresponding coefficients in the lower triangle as their complex conjugates. From the improper APT guarantee, we have
@@
$\tr_{1}(J(\Phi_{x,y})) =tr$$=\tr_{1}\left(\sum_{a,b}\sigma_{x}|a\rangle\langle b|\sigma_{y}\otimes|a\rangle\langle b|\right)$ (5.193)
@@
This transformation is well-known, first noted in [goemans2001approximation], and can be used to transform any complex semidefinite programming problem into its real analogue. It can be shown that this transformation preserves the properties of the original operator. For example, a Hermitian operator is mapped to a symmetric one and a Hermitian positive semi-definite operator is mapped to a symmetric positive semi-definite one. Using this transformation, we see that it suffices to solve the following convex optimization problem over reals:
@@
$C_{1}=\{x\in\mathbb{R}^{4D^{2}}\mathrel{\mathop{\ordinarycolon}}\mathsf{vec}^{-1}(x)\in S_{2D}^{+}\},\qquad C_{2}=\{x\in\mathbb{R}^{4D^{2}}\mathrel{\mathop{\ordinarycolon}}\tilde{A}x=\tilde{b}\}$, (5.211)
@@
Note that projecting onto either $C_{1}$ or $C_{2}$ individually is easy (running in time $\mathrm{poly}(D)$) since they are simply the PSD cone and a polyhedral set (see, e.g., [henrion2012projection]). Note that the runtime being in terms of $D$ is not a problem for our applications, as we discuss at the end of this section. Thus, the problem is as follows: how can we use projections onto $C_{1}$ and $C_{2}$ individually to project onto $C_{1}\cap C_{2}$? This problem is well-studied in convex optimization, where researchers have developed methods such as alternating projections [von1949rings, von1950functional] and Dykstra’s algorithm [dykstra1983algorithm]. As the name suggests, at a high level, these algorithms work by projecting the input onto $C_{1}$ and $C_{2}$ in turn. Classical results in optimization theory show that Dykstra’s algorithm converges to the desired projection [dykstra1983algorithm, boyle1986method, bauschke1994dykstra, han1988successive, bauschke2000dykstras], and its variants are also widely studied [bauschke2020dykstra, deutsch1994rate, gaffke1989cyclic, hundal1997two, pang2015set]; however, researchers have only recently rigorously analyzed its convergence rate for non-polyhedral sets [wang2024convergence]. Convergence rates have previously been analyzed for the basic alternating projections algorithm [deutsch2001best, bauschke1993convergence, lewis2009local] and Dykstra’s algorithm for projection onto the intersection of polyhedral sets [deutsch1994rate, tseng2009coordinate, luo1993error]. However, the former only guarantees convergence to a point in $C_{1}\cap C_{2}$, rather than the closest one, which would not help for our purposes since we require an error bound. Also, the latter does not apply to our setting, as $C_{1}=S_{2D}^{+}$ is clearly not polyhedral. In our setting, using Dykstra’s algorithm, one would hope for linear convergence, as this leads to a computationally efficient (in the size of the input matrix) projection algorithm. In fact, [wang2024convergence] shows that Dykstra’s algorithm achieves a linear convergence rate under certain conditions. Namely, they prove the following theorem.
Theorem 10 (Theorem 5.3(i) in [wang2024convergence]; Simplified).
@@
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\mathinner{\!\left\lVert x-x_{0}\right\rVert}_{2}^{2}\qquad\text{s.t.}\qquad x\in\bigcap_{i=1}^{\ell}C_{i},$ (5.212)
@@
    $0\in x^{*}-x_{0}+\mathrm{ri}\left(\partial\left(\sum_{i=1}^{\ell}\delta_{C_{i}}\right)(x^{*})\right)$, where $x^{*}$ is the optimal projected point, $\partial$ denotes the subdifferential, and $\delta_{C_{i}}$ is the indicator function of the set $C_{i}$ (0 for inputs inside $C_{i}$ and $\infty$ otherwise).
@@
Note that [wang2024convergence] considers a more general problem, in which they wish to project onto a point such that a linear transformation of this point lies in the intersection of the convex sets. For our setting, we can take this transformation to be the identity, in which case the algorithm from [wang2024convergence] reduces to Dykstra’s algorithm. Thus, this theorem would give the convergence result we need if its conditions hold.
Now, let us discuss each of the conditions. First, for Condition 1, the exact definition of $C^{1,\alpha}$-cone reducibility, for $\alpha\in(0,1]$, is outside of the scope of this work, and we refer to Definition 4.1 in [wang2024convergence] for the definition. Importantly, this is an extension of the more common notion of $C^{2}$-cone reducibility [shapiro2003sensitivity, bonnans2013perturbation]. Specifically,  [wang2024convergence] states that any $C^{2}$-cone reducible set is also $C^{1,\alpha}$-cone reducible for all $\alpha\in(0,1]$. Moreover, [shapiro2003sensitivity] shows that all polyhedrons and the cone of positive semidefinite matrices are both $C^{2}$-cone reducible. Hence, our convex sets $C_{1}$ and $C_{2}$ are both $C^{2}$-cone reducible and hence $C^{1,\alpha}$-cone reducible for all $\alpha\in(0,1]$. In particular, $C_{1}$ and $C_{2}$ are both $C^{1,1}$-cone reducible, satisfying Condition 1.
@@
Also, $\partial$ denotes the subdifferential. For a function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{4D^{2}}\to\mathbb{R}$, the subdifferential is defined as
$\partial f(x)\triangleq\{z\in\mathbb{R}^{4D^{2}}\mathrel{\mathop{\ordinarycolon}}f(y)\geq f(x)+(y-x)^{\intercal}z\qquad\forall y\in\mathbb{R}^{4D^{2}}\}.$ (5.214)
@@
$\mathrm{ri}(S)\triangleq\{x\in S\mathrel{\mathop{\ordinarycolon}}\forall y\in S,\;\exists\lambda>1\text{ s.t. }\lambda x+(1-\lambda)y\in S\}.$ (5.215)
Because of Condition 2, standard results in subdifferential calculus (see, e.g., [nesterov2018lectures]) state that
@@
$N_{S}(y)\triangleq\{a\in\mathbb{R}^{4D^{2}}\mathrel{\mathop{\ordinarycolon}}a^{\intercal}y\geq a^{\intercal}x,\qquad\forall x\in S\}.$ (5.217)